Lab: Trend Projection is a time-series forecasting models that examines historical data. The quantitative goal is to fit the data trend line to a series of historical data points. The qualitative goal is to use the data as a means to project medium to long-term forecast plans. In Industry, this is the method to reduce the SSE, Sum of Squares Errors.

Given the following data, create a Trend Projection with Least Squares Analysis. Compute the slop and intercept of the line. Use the Slope and Intercept data to project demand for 2018. Hint: You need to change the Data to reflect the more recent years:

Boston Data Servers Inc.: Demand Forecast

Forecast Year

Data Server Demand

2011

91

2012

101

2013

85

2014

99

2015

147

2016

180

2017

137

5. To evaluate the model, we compare the historical demand and the trend line. Please comment on your personal findings.
Answer #1

PERIOD

DEMAND

X

Y

XY

X^2

1

91

1

91

91

1

2

101

2

101

202

4

3

85

3

85

255

9

4

99

4

99

396

16

5

147

5

147

735

25

6

180

6

180

1080

36

7

137

7

137

959

49

SIGMA

28

840

3718

140


INTERCEPT(B0) = (SIGMA(Y) * SIGMA(X^2) - SIGMA(X) * SIGMA(XY)) / (N * SIGMA(X^2) - SIGMA(X)^2)

INTERCEPT = (840 * 140) - (28 * 3718) / ((7 * 140) - 28^2) = 68.86

SLOPE(B1) = ((N * SIGMA(XY)) - (SIGMA(X) * SIGMA(Y))) - (N * SIGMA(X^2) - SIGMA(X)^2)

SLOPE = ((7 * 3718) - (28 * 840) / ((7 * 140) - 28^2) = 12.79

EQUATION

Y = INTERCEPT + SLOPE(X)

Y = 68.86 + (12.79 * (X))


FOR X = 8 FORECAST 2018 = 68.86 + (12.79 * 8) = 171.18

2. WITH THE SLOPE VALUE, WE CAN SAY THAT THE SERVER DEMAND INCREASES BY 12.79 PER YEAR.